OpenStudy (anonymous):

hich expression is equivalent to this one? k • 6 + k • s A. 6(k + s) B. s(6 + k) C. k(6 + s) D. k · (6 + k) · s

OpenStudy (anonymous):

@TheMysteriousRainbowPegasus

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Lol, copy/paste fail "hich expression is equivalent to this one?" Which one do you think it is?

OpenStudy (anonymous):

lol ok i think it is C

OpenStudy (anonymous):

Why do you think this answer is correct? Give your reasoning.

OpenStudy (anonymous):

well it is because you are doing the equation and you find the answer and mach it up

OpenStudy (anonymous):

with other

OpenStudy (anonymous):

Oops, sorry... it's asking what is pretty much the same question is, but with different things... if that makes sense? For this question, the answer would be D. It has everything the original question does, but with the marking in a different place. Understand? Sorry, I'm not the best at explaining things.

OpenStudy (anonymous):

it is ok i understand and man it is raining hard here

OpenStudy (anonymous):

so it is D

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have 2 more after this.

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

What is the solution to the equation? 7(5 + 9) = 7 • x + 7 • 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (anonymous):

Okay, do you understand what it's asking?

OpenStudy (anonymous):

no not this one maby solve it

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Okay, hold on.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It's definitely A.

OpenStudy (anonymous):

ok and Use the distributive property to rewrite the expression and solve. 9 72 = 9 + 9 2 =

OpenStudy (anonymous):

I'm not good at properties.. sorry. :c

OpenStudy (anonymous):

I'll find someone to tag.

OpenStudy (anonymous):

ok do you know who?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'll find someone to tag.

OpenStudy (anonymous):

@Preetha